OpenStudy (anonymous):

The hypotenuse of a right triangle is 17. The lengths of the two legs are whole numbers. What are they?

OpenStudy (anonymous):

15 and 8

OpenStudy (anonymous):

\[17^2 = 289\] \[15^2 +8^2 = 225 + 64 = 289\] this followa that: \[17^2 = 15^2 + 8^2\] Hence remaining two sides are 15 and 8

OpenStudy (anonymous):

Thank you

OpenStudy (anonymous):

One leg of a right triangle has length 7. The lengths of the other two sides are whole numbers. The length of the other leg is BLANK and the length of the hypotenuse is BLANK

OpenStudy (anonymous):

HELP

OpenStudy (anonymous):

 is this what you mean?

OpenStudy (anonymous):

or is the hypotnuse 7?

OpenStudy (anonymous):

The hypotenuse is blank, I'm trying to figure out the hypotnuse. the (7) Is the length of a leg

OpenStudy (anonymous):

hello

OpenStudy (anonymous):

hello

OpenStudy (anonymous):

sorry, i had to go somewhere. a^2+b^2=c^2 right? so 7^2+b^2=c^2 49+b^2=c^2 i'll square root everything. so 7+b=c, or 7=c-b here we can just sub in random numbers, as long as the hypotenuse is the biggest and as long as the make perfect square roots. so i will try b=3. if you sub it in, c becomes 10, because 7+(3)=c 10=c. it can work with many other numbers, as long as c is the longest.